OpenStudy (anonymous):

the length of a rectangle is three times its width. if the perimeter of the rectangle is 48 cm, find its area

OpenStudy (anonymous):

ok, so L=3w so your equation should be 2(3w)+2w=48

OpenStudy (anonymous):

w=6?

OpenStudy (anonymous):

and L=18?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

and the area is 108

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf length = l\qquad width = w\\ length = 3w\\ \quad \\ \quad \\ perimeter = 48 = 2l+2w = 2(3w)+2w\qquad thus\\ \quad \\ 48 = 2(3w)+2w\)

OpenStudy (anonymous):

Yea. :)